Express the following in rectangular and polar form, if


Z1 = 3+ 4i


Z2= 2+3i


1. Z1*Z2


2. Z1-Z2


3. Z1/Z2


4. |Z1|


5. |Z2|



(2)if Z1=50<30°and Z2=30<60°find in rectangular


form the following


1. |Z1|


2. |Z2|


3. |Z1|-|Z2|


4. Z1*Z2


5. |Z2-Z1|


6. |Z2|/|Z1|

1
Expert's answer
2021-11-11T12:28:55-0500

(1)

1.

"z_1\cdot z_2=(3+ 4i)(2+ 3i)=6+9i+8i-12"

"=-6+17i"

"|z_1\cdot z_2|=\sqrt{(-6)^2+17^2}=5\sqrt{17}"

"\tan \theta=\dfrac{17}{-6}, 90\degree<\theta<180\degree"

"z_1\cdot z_2=5\sqrt{17}\angle\big(180\degree-\tan^{-1}\dfrac{17}{6}\big)"

2.

"z_1-z_2=(3+ 4i)-(2+ 3i)=1+i"

"|z_1- z_2|=\sqrt{1^2+1^2}=\sqrt{2}"

"\tan \theta=\dfrac{1}{1}=1, 0\degree<\theta<90\degree"

"z_1-z_2=\sqrt{2}\angle45\degree"

3.

"z_1/ z_2=(3+ 4i)/(2+ 3i)=\dfrac{(3+4i)(2-3i)}{2^2+3^2}"

"=\dfrac{6-9i+8i+12}{13}=\dfrac{18}{13}-\dfrac{1}{13}i"

"|z_1/ z_2|=\sqrt{(\dfrac{18}{13})^2+(-\dfrac{1}{13})^2}=\dfrac{5\sqrt{13}}{13}"

"\tan \theta=\dfrac{-\dfrac{1}{13}}{\dfrac{18}{13}}=-\dfrac{1}{18}, 270\degree<\theta<360\degree"

"z_1/ z_2=\dfrac{5\sqrt{13}}{13}\angle\big(360\degree-\tan^{-1}\dfrac{1}{18}\big)"

4.


"|z_1|=\sqrt{3^2+4^2}=5"

"|z_1|=5+0i"

"|z_1|=5\angle0\degree"

5.


"|z_2|=\sqrt{2^2+3^2}=\sqrt{13}"

"|z_2|=\sqrt{13}+0i"

"|z_2|=\sqrt{13}\angle0\degree"

(2)

1.


"|z_1|=50+0i"

2.


"|z_2|=30+0i"



3.


"|z_1|-|z_2|=50-30=20"


"|z_1|-|z_2|=20+0i"



4.


"z_1\cdot z_2=50\cdot 30(\cos (30\degree+60\degree)+i\sin (30\degree+60\degree))"




"=1500(0+i(1))"


"z_1\cdot z_2=0+1500i"

5.


"z_1=50(\cos 30\degree+i\sin 30\degree)=25\sqrt{3}+25i"

"z_2=30(\cos 60\degree+i\sin 60\degree)=15+15\sqrt{3}i"


"z_2-z_1=(15-25\sqrt{3})+(15\sqrt{3}-25)i"

"|z_2-z_1|=\sqrt{(15-25\sqrt{3})^2+(15\sqrt{3}-25)^2}"

"=\sqrt{225-750\sqrt{3}+1875+675-750\sqrt{3}+625}"

"=10\sqrt{34-15\sqrt{3}}"

"|z_2-z_1|=10\sqrt{34-15\sqrt{3}}+0i"

6.


"\dfrac{|z_2|}{|z_1|}=\dfrac{30}{50}=\dfrac{3}{5}"


"|z_2|/|z_1|=\dfrac{3}{5}+0i"